OpenStudy (anonymous):

(2x+3)^2=5 solve by using the square root property

OpenStudy (anonymous):

(2x+3)^2 = 5 (2x+3) =\[\sqrt{5}\] 2x = 3+\[\sqrt{5}\] x = 3+\[\sqrt{5}\] / 2

OpenStudy (anonymous):

2x + 3 = +- sqrt5 x = (sqrt5 - 3) / 2 or (-sqrt5 - 3) / 2

OpenStudy (anonymous):

= -0.38 or -2.62